Soft beverages contain large amounts of carbonic acid (H2CO3, a diprotic acid) and phosphoric acid (H3PO4, a triprotic acid). The total acid concentration in a given beverage can be determined by performing an acid-base titration with sodium hydroxide. The label on a 500.0mL soft drink bottle indicates that the concentrations are: H2CO3 - 1.33x10^-4 mol/L and H3PO4 - 3.85x10^-3 mol/L. Based on these concentrations, what volume of sodium hydroxide solution(0.10M) would be required to completely react with all the acidic species present in this bottle?
1
Expert's answer
2017-09-18T08:39:06-0400
n(H₂CO₃)=1.33*0.5⁻⁵Mol
n(H₃PO₄)=3.85*10⁻³*0.5=1.925*10⁻³Mol
H₂CO₃+2NaOH->Na₂CO₃+2H₂O
n1(NaOH)=6.65*10⁻⁵*2=1.33*10⁻⁴Mol
H₃PO₄+3NaOH->Na₃PO₄+3H₂O
n2(NaOH)=1.925*10⁻³*3=5.775*10⁻³Mol
n(NaOH)1.33*10⁻⁴+5.775*10⁻³=5.908*10⁻³Mol
V(NaOH)=5.908*10⁻³:0.1=59.08ml